If a solution of urea (CO(NH2)2) has a mole fraction of 0.815, what is its molality? (Use 60.063 g/mole as the molar mass of urea for any equations that require the use of the molar mass of urea.)(8 Points)
Answer #1

Here the mole fraction of urea is given as 0.815

So mole fraction of water is going to be 0.185

Using this value we can calculate mass of both the constituents

Mass of urea = molar mass×mol fraction

= 60.063× 0.815

=48.95 gm

Similarly mass of water = 18× 0.185

= 3.33gm

To calculate molarity

Molality=

= 0.815/3.33

=0.244

so the solution will have 0.244 molality

i hope this helps if you have any query or want more detailed explanation feel free to ask in the comment section below.